OpenStudy (anonymous):

how do i get the volume of the right circular cone generated by rotating the line y =2/3 x about the y-axis between y = 0 and y = 3.

OpenStudy (campbell_st):

well it's and inverted cone  so the volume is \[V = \pi \int\limits_{0}^{3} x^2 dy\] so take the equation y = 2/3x and make x the subject so x = 3/2y now square it so you get \[x^2 = \frac{9y^2}{4}\] now substitute and the integral is \[V = \pi \int\limits_{y = 0}^{y = 3} \frac{9y^2}{4} dy\] so just integrate with respect to y. hope it helps